# Mitigate Position Bias in Large Language Models via Scaling a Single Dimension

### v2

@@ 3.2 Identifying Positional Hidden States
We have defined positional hidden states in Definition 2.1. However, the original values of hidden states may not strictly satisfy monotonicity. After curve fitting, we can identify dozens or hundreds of dimensions that exhibit various degrees of relevance to positional information. Thus, the first step of our method is to find the dimension that best fits the properties of positional hidden states.
@@
To efficiently search for the positional hidden states from the LLMs’ hidden states set, we leverage the characteristics of positional hidden states defined in Section 2.3 and propose a heuristic positional hidden search algorithm. As shown in Algorithm 1, the search process consists of the following two steps: 1) Identify the top-$k$ dimensions $\rho$ in the hidden states that are monotonic in more than $\varepsilon$ layers and are as smooth as possible. Here $c_{t}$ is the number of layers where $h_{t}(p)$ is monotonic, and $g_{t}$ is the smooth score of $h_{t}(p)$. Equ.(3) is the smoothness formula. 2) Use a small validation dataset (or called calibration dataset) $\mathcal{D}_{\text{val}}=\{\bm{x},\bm{y}\}$ to evaluate the impact of scaling these positional hidden states respectively and select the positional hidden states $h_{\overline{t}}$ that can lead to the minimal loss $\mathcal{L}_{\overline{t}}$.
$\text{Smooth}(h_{t})=\int|h^{\prime\prime}_{t}(p)|^{2}$ (3)
As for selecting the best scale factor, we take 0.5, 0, -0.5, and -1 to respectively experiment on the validation set, obtain the validation loss, and then select the scaling factor with the lowest loss.
3.3 Scaling the Positional Hidden States
@@
To minimize the impact of this modification on the semantics of LLMs, we propose a method scaling the positional hidden states only affecting the last token, as shown in Figure 4. Specifically, for the tokens preceding the last token, the attention calculation remains the same as the original. For the last token’s attention computation of a sequence of length $l$, we obtain the modified query state $\bm{\overline{q}}_{l}$ (of the $l$-th token, i.e. the last token) and key states $\bm{\overline{K}}$ (of all the tokens) by scaling the positional hidden states. That is,
$\bm{\overline{q}}_{l}=\mathcal{P}(W^{Q}f(\bm{h}(l),p,s),l),\quad\bm{\overline{K}}=\mathcal{P}(W^{K}f(\bm{h},p,s),[1,2,...,l])$ (4)
@@
Here $f(\bm{h},p,s)$ means the $p$-th dimension of $\bm{h}$ is scaled by the factor $s$. Therefore, the corresponding attention calculation is as follows:
@@
where ${\bm{z}}$ is the attention output. We implement our method using FlashAttention (Dao, 2024) with minimal overhead. After calculating the combined attention weights, the remaining computations remain the same as in the original method. As shown in Appendix A.4, our approach results in only a slight increase in latency.
@@ 4.1 Setup
We apply our method to a wide range of state-of-the-art open-source LLMs, including: 1) RoPE (Chen et al., 2023) models: LLaMA-2 (7B, 13B) (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma-7B (Team et al., 2024), Qwen1.5-7B (Bai et al., 2023); 2) Context window extended models: Vicuna (7B, 13B) (Chiang et al., 2023); 3) Alibi (Press et al., 2022) models: MPT-30B (Team, 2023). All the models we use are instruction-tuned versions.
And we evaluate the performance across three aspects: 1) Position-bias-related tests on NaturalQuestion multi-document QA (Liu et al., 2024c) and KV retrieval (Liu et al., 2024c) with ground-truth at different positions in the prompt. The NaturalQuestion task includes 20 documents with a prompt length of about 2.3k tokens, while the KV retrieval task includes 140 KV pairs with an average length of about 10k tokens. 2) General long-context benchmark on LongBench (Bai et al., 2024), including multi-document QA, single-document QA, summarization, few-shot learning, synthetic tasks, and code completion, totaling 16 tasks with an average length of 37k tokens. 3) Position-sensitive tasks on timeline reordering in LooGLE (Li et al., 2024a), with an average length of 10k tokens. For prompts that exceed the context windows of LLMs, we follow LongBench’s approach by truncating from the middle and retaining the head and tail of the prompt to fit within the context windows. We use the provided metrics and scripts from the following benchmarks for evaluation.
@@
In this paper, we implement our approach using PyTorch, HuggingFace Transformers, and FlashAttention (Dao, 2024) in an A100 GPU. To ensure stable and reproducible results, we use greedy decoding in all experiments. For the search part, we set the top-$k$ size of positional hidden states to 10 and $\varepsilon$ to $L/4$, where $L$ is the number of layers. The validation set is a synthetic KV retrieval dataset consisting of 100 examples, which do not overlap with the test set. The search process takes approximately 10 minutes. For the scaling part, we only modify the intermediate layers of the model to minimize the negative impact on performance. The details of the scaling dimensions, layer ranges, and factors are shown in Table 5. More details are provided in Appendix A.
@@
We include two training-free positional bias mitigation methods as baselines: (i) Original, the unmodified LLM results with the ground truth at different positions in the prompt. (ii) w/ Ms-PoE (Zhang et al., 2024), a head-aware position embedding scaling method to mitigate position bias. Following the original settings, we apply scaling coefficients of 1.2 to 1.8 starting from the 3rd layer.
@@ 4.2 Main Results
Table 1: Performance of different methods with different models on NaturalQuestions (20 docs) (Liu et al., 2024c) and KV retrieval (140 KV pairs) (Liu et al., 2024c) dataset.
@@
Tables 1 and 2 present the performance of various methods in different benchmarks. Several observations and conclusions can be drawn: 1) Our method consistently improves overall performance at different positions, with increases of up to 9.3%, 15.2%, and 4.7% in NQ, KV retrieval, and LongBench, respectively, except for LLaMA-2-13B in KV retrieval. Additionally, compared to the SoTA method Ms-PoE, our method shows significant improvements of up to 6.3%, 97.5%, and 14% in NQ, KV retrieval, and LongBench. The poor performance of Ms-PoE in KV retrieval can be attributed to the interpolation causing information loss. 2) Our method effectively enhances LLMs’ understanding of information located in the middle and latter parts of the prompt. For key information at the beginning of the prompt, performance is comparable to baselines. Considering only the average performance of the last four positions, our method’s improvements over the original increase to 11.3% and 16.8% in NQ and KV retrieval, respectively, and over Ms-PoE increase to 8.7% and 97.5% in NQ and KV retrieval, respectively. 3) Our approach is effective not only for RoPE models but also for context window extended models like Vicuna-16K, which already readjust RoPE (Chen et al., 2023). Additionally, our method can be adapted to different position embeddings, such as Alibi (Press et al., 2022) models like MPT, resulting in improvements of 2.6%, 15.2%, and 1.2% in NQ, KV retrieval, and LongBench, respectively. 4) Our method demonstrated varying degrees of improvement across different tasks, with the most significant increases being 1.5% in few-shot learning tasks, 3.4% in code tasks, 4% in synthetic tasks, 9.2% in single document QA tasks, and 1.9% in multi-document QA tasks. In summarization tasks, performance was nearly on par with the original results. While our method did not significantly improve the average scores overall, it at least demonstrates that it can mitigate position bias without impairing the model’s original capability to handle long context tasks.
@@ 4.3 Analysis
As shown in Table 1, there is an phenomenon that our method mainly benefits when the key information is not at the beginning, but can often decrease performance if the model performs significantly better when the key information is at the beginning. It reveals a possible fact that the positional hidden may be an important factor causing the model to miss the rear parts of the context while focus too much to the beginning parts. Therefore, scaling such dimension can shift the model’s attention from being too focused at the beginning to a more balanced distribution. We validated the above points by testing different scale factors, as shown in Figure 5.
Scale Factor
(a) Scale factor v.s. Attention
(b) Scale factor v.s. Performance
Figure 5: Attention distribution and performance when scaling dimension 2393 of Vicuna-7b-v1.5-16k with different scale factors on KV retrieval (Liu et al., 2024c) of 100 KV pairs.
The scaling factor directly controls the degree and direction of the impact of positional hidden states on position bias. As shown in Figure 5, a positive scaling factor causes the model to focus more on the beginning, while a negative factor shifts the focus towards the end. A factor between 0.5 and -1 leads to the most balanced attention distribution, where accuracy also peaks. These results demonstrate that scaling positional hidden states can influence LLMs’ tendency to focus on the beginning, and by adjusting the coefficients, this bias can be effectively mitigated.
Ablation Study
Method LLaMA-2-7b Vicuna-13b Gemma-7b Mistral-7b Qwn1.5-7b
Original 31.3 50.2 28.0 59.5 56.8
Ours 40.6 52.7 34.6 60.9 58.3
w/o monotonicity 40.6 51.8 34.6 60.9 58.3
w/o smoothness 40.6 52.7 27.8 60.9 58.3
w/o validation set 30.1 51.8 26.5 60.9 58.3
w/ scale 2 dimensions 37.2 50.8 31.7 60.1 57.2
w/ modify last 16 tokens 41.6 51.5 34.6 59.7 58.1
w/ modify all tokens 44.0 50.8 31.7 59.5 57.4
Table 3: Average performance of different ground-truth positions using different methods on NaturalQuestions multi-document QA dataset (20 docs) Liu et al. (2024c).
To evaluate the contributions of different components in our method, we introduce the following sets for the ablation study: (1) Ours w/o monotonicity, w/o smoothness, and w/o validation set, which adjust the search algorithm by not considering these three indicators, respectively (details in Appendix A.2). (2) Ours w/ scale 2 dimensions, which modifies the top-2 positional hidden states simultaneously. (3) Ours w/ modify last 16 tokens and w/ modify all tokens, which adjust the range of tokens affected by the scaling operation in Equ.(5).
Table 3 shows the ablation results. It can be seen that without filtering by monotonicity or smoothness, performance may decline, and removing the validation set results in more decline in model performance. When the range of tokens or dimensions affected by scaling is expanded, most models experience varying degrees of performance loss. Considering these factors, we choose to modify only the last token and the top-1 positional dimension to achieve the best performance.
Side Effects
We utilized the MMLU dataset (Hendrycks et al., 2020), which assesses general capabilities, and the timeline-reorder dataset (Li et al., 2024a), which is a task sensitive to positional information, to evaluate whether our approach adversely affects the original abilities of the LLM.
Attention v.s. Performance
(a) Attention
(b) Attention w/ modification
(c) Accuracy
(d) Accuracy w/ modification
Figure 6: Distribution of attention weight and accuracy as the ground-truth KV is placed at different positions in the prompt. (b) and (d) are situations when the attention on the 25th KV pair is modified.
As shown in Figure 6, when we manually double the attention weights of the key information (in this case, the 25th KV pair, as illustrated in Figure 6b) during the model’s forward pass on the KV retrieval task, the retrieval accuracy for the 25th KV improves, while the accuracy for other parts remains largely unchanged (Figure 6d). This demonstrates that the attention weights for key information are positively correlated with retrieval accuracy.
Does this Method Compromise the Ability to Perceive Positional Information?
@@
Table 4: Performance of difference models on MMLU and the timeline reorder task.
To demonstrate that our method does not harm the model’s performance on general or position-sensitive tasks, despite eliminating some positional information, we tested it on two datasets: the MMLU benchmark (Hendrycks et al., 2020) and the timeline reorder task from LooGLE (Li et al., 2024a), which involves arranging events chronologically across extensive text. As shown in Table 4, our method does not impair performance on position-sensitive tasks, indicating that the positional information we remove may not be essential for the model’s functioning.
5 Related Works
Long-Context LLMs
Recent research has focused on expanding the context window size of LLMs through four main approaches: 1) Staged pre-training (Nijkamp et al., 2023; Fu et al., 2024), which gradually increases the context window size during training; 2) Modifying or interpolating position embeddings (Press et al., 2022; Chen et al., 2023; Peng et al., 2024; Ding et al., 2024); 3) Using external memory modules for context storage (Bertsch et al., 2023; Tworkowski et al., 2023); 4) Distributed computation across devices (Liu et al., 2024b). While these methods address context expansion, their impact on positional bias in downstream tasks has not been thoroughly explored.
@@ 5 Related Works
Despite explicit positional encoding methods like RoPE (Su et al., 2024) and Alibi (Press et al., 2022), LLMs often exhibit position bias, such as the "lost in the middle" phenomenon (Liu et al., 2024c; Kamradt, 2023). Recent efforts to mitigate this bias fall into several categories: 1) RoPE-based methods: These approaches modify the RoPE computation process to alleviate long-distance information decay, including Attention Bucket (Chen et al., 2024), which uses an ensemble of multiple RoPE bases to mitigate position bias, and Ms-PoE (Zhang et al., 2024), which dynamically interpolates with a small coefficient for different heads. 2) SFT-based methods (Junqing et al., 2023; Yu, 2023; An et al., 2024): These methods construct data with more diverse key information distributions or employ system2think SFT tasks to mitigate position bias. They require further training of the model. 3) Attention mask-based methods (He et al., 2024): These methods modify attention mechanisms, including Attention Transition (Gao et al., 2023), which redirects attention to significant parts of the context and Stable Mask (Yin et al., 2024), which introduces pseudo attention into the causal mask, ensuring stable attention distribution when facing lengthy texts. 4) Prompt-based methods (Jiang et al., 2024; Peysakhovich & Lerer, 2023): These methods introduce an external module to reorder or compress information in the prompt, thereby mitigating position bias.
6 Conclusion
This paper proposes a method for scaling positional hidden states to mitigate position bias issue in LLMs. Specifically, the study first confirms that attention weights manifest position bias within transformers. Additionally, experiments demonstrate that, besides position embeddings, the causal mask also contributes to position bias, which is transmitted to other modules through the hidden states containing absolute positional information, termed as positional hidden states. Based on this, we introduce a prior-based positional hidden search algorithm and mitigate the model’s position bias by scaling the positional hidden states searched. Testing eight open-source models with different position embeddings on tasks such as NaturalQuestions Multi-document QA, KV Retrieval, and LongBench, the results show that our method effectively reduces position bias and improves model performance.
Appendix A Experiment Details
A.1 Datasets Details

### v3

@@ 3.2 Identifying Positional Hidden States
We have defined positional hidden states in Definition 2.1, but it is just an ideal situation. In practice, because the hidden states is mainly determined by specific input text, the original values of them will not strictly satisfy monotonicity, which means we cannot rely on strict monotonicity to select the positional channels, but whether it roughly conforms to monotonicity. Thus, we use least square polynomial fit to approximate the values in the channel. Moreover, hidden states of different layers also vary. So we consider this channel as roughly monotonic if the fitted curve conforms to monotonicity in more than a quarter of all the layers of the model.
Using curve fitting, we can usually identify dozens or hundreds of channels that exhibit various degrees of monotonicity. However, some of them are very close to the ideal monotonic curve, while others fluctuate violently although they are roughly monotonic. So we use another indicator, smoothness, to evaluate if how close they are to the ideal monotonic change. The smoothness score is calculated by $\int|h^{\prime\prime}_{t}(p)|^{2}$, where $h^{\prime\prime}(p)$ is the second derivative (or difference) of $h(p)$, and a smaller score means smoother. Only when $h_{t}$ is considered roughly monotonic, its smoothness score will be calculated, and averaged across layers. Then we only maintain the channels with the Top-$K$ smoothness score. $K$ is set to 10 in our default setting.
Finally, we evaluate the average loss across $K$ channels using a 100-sample calibration dataset for KV retrieval. The channel with the lowest loss is chosen for scaling.
To determine the optimal scale factor, we perform a grid search over {0.5, 0, -0.5, -1}, selecting the factor with the lowest loss.
@@
We organize our search algorithm in Algorithm 1, the search process consists of the following two steps: 1) Identify all the channels $\rho$ of the hidden states that are roughly monotonic in more than $\varepsilon$ layers, and select the top-$K$ channels with the lowest smoothness score. Here $c_{t}$ is the number of layers where $h_{t}(p)$ is monotonic, and $g_{t}$ is the smooth score of $h_{t}(p)$. 2) Use a small calibration dataset $\mathcal{D}_{\text{val}}=\{\bm{x},\bm{y}\}$ to evaluate the impact of scaling these positional hidden states respectively and select the channel $h_{\overline{t}}$ that can lead to the minimal loss $\mathcal{L}_{\overline{t}}$.
3.3 Scaling Positional Hidden States
In our early experiments we find simply scaling the positional channel of the hidden states in every layer or in every module of the model can lead to unstable performance of the model. Therefore, to minimize the side-effect of scaling positional hidden states, our proposed method is finely designed to scale the positional hidden states only affecting the attention weights from last token of the sequence, as shown in Figure 5.
Due to the inherent lack of interpretability in the working mechanisms of LLMs, our design is mainly based on the experience of trial and error experiments. So, we use some ablation experiments, shown in Appendix E, to explain why we scale only 1 channel but not more, modify the last token’s attention but not all tokens, and use 3 indicators to search the optimal channel.
@@ 3.3 Scaling Positional Hidden States
Specifically, in our method, for the tokens preceding the last token, the attention calculation remains the same as the original. In a sequence of length $l$, for the last token’s attention computation, we obtain the modified query state $\bm{\overline{q}}_{l}$ (of the $l$-th token, i.e. the last token) and key states $\bm{\overline{K}}$ (of all the tokens) by scaling the positional hidden states. That is,
$\begin{split}&\bm{\overline{q}}_{l}=\mathcal{P}(W^{Q}f(\bm{h}(l),p,s),l)\\ &\bm{\overline{K}}=\mathcal{P}(W^{K}f(\bm{h},p,s),[1,2,...,l])\end{split}$ (2)
@@
Here $f(\bm{h},p,s)$ means the $p$-th channel of $\bm{h}$ is scaled by the factor $s$. Therefore, the combined attention calculation is as follows:
@@
where ${\bm{z}}$ is the attention output.
Except calculating the combined attention weights, the other modules remain the same as those in the original method. We implement our method using FlashAttention-2 (Dao, 2024) with minimal overhead, so our approach results in only a slight increase in latency, as shown in Appendix C.4,
@@
Moreover, we find modifying the hidden states in the initial layers (especially the 1st or 2nd layer) will also make the model’s output more unstable, as well as the last layers. Thus we only apply our scaling to layers in the middle (they are more likely to be retrieval-related layers), for example, the 10th to 25th layers with total 32 layers. The specific layer selection is also based on engineering experience.
@@ 4.1 Setup
We apply our method to a wide range of popular open-source LLMs, including: 1) RoPE (Chen et al., 2023) models: LLaMA-2 (7B, 13B) (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma-7B (Team et al., 2024), Qwen1.5-7B (Bai et al., 2023); 2) Context window extended models: Vicuna-16k (7B, 13B) (Chiang et al., 2023); 3) Alibi (Press et al., 2022) models: MPT-30B (Team, 2023). All the models are instruction-tuned versions.
And we evaluate the performance in 2 aspects: 1) Position-bias-related tests: “lost in the middle” benchmark, including NaturalQuestion multi-document QA (Liu et al., 2024) and KV retrieval (Liu et al., 2024) with the Gold document or KV at different positions in the context. The multi-document QA task includes 20 documents with a prompt length of about 2.3k tokens, while the KV retrieval task includes 140 KV pairs with an average length of about 10k tokens. 2) General long-context multi-task benchmark: LongBench (Bai et al., 2024), including multi-document QA, single-document QA, summarization, few-shot learning, synthetic tasks, and code completion, totaling 16 tasks with an average length of 37k tokens.
For prompts that exceed the context windows of LLMs, we follow LongBench’s approach by truncating from the middle and retaining the head and tail of the prompt to fit within the context windows. We use the metrics and scripts provided by the benchmark’s github repository for evaluation. More details about the benchmarks are in Appendix B.
@@
We implement our approach using PyTorch and HuggingFace Transformers in an A100 GPU. To ensure stable and reproducible results, we use greedy decoding in all generation experiments.
In the searching algorithm, we set the top-$K$ to top-10 and $\varepsilon$ to $L/4$, where $L$ is the number of total layers. The search process takes approximately 10 minutes. The details of the scaling channels, layers, and factors are shown in Appendix C.3.
@@
Besides the original models, we include another training-free positional bias mitigation method as the baseline, which is signified by w/ Ms-PoE (Zhang et al., 2024). It is a head-wise position embedding scaling method, and we follow its default settings to apply scaling coefficients of 1.2 to 1.8 starting from the 3rd layer.
@@ 4.2 Main Results
Table 1: Performance of different methods with different models on NaturalQuestions (20 docs) (Liu et al., 2024) and KV retrieval (140 KV pairs) (Liu et al., 2024) dataset. “/ ” signifies “not applicable”.
@@
From the evaluation results of “lost in the middle” benchmark in Table 1, several conclusions can be drawn: 1) Our method better improves overall performance at most positions, with the average improvement of up to 9.3%, 15.2% in NQ and KV retrieval, respectively, except for LLaMA-2-13B in KV retrieval. 2) Our method effectively enhances LLMs’ exploitation of information located in the middle and rear parts of the prompt. When key information is at the beginning of the prompt, performance is sometimes increased or decreased. Considering only the average performance of the last four positions, our method’s improvements over the original increase to 11.3% and 16.8% in NQ and KV retrieval, respectively, much higher than MsPoE. 3) Our method demonstrates better generalization performance, showing improvement on nearly all types of models, regardless of RoPE models, context window extended models like Vicuna-16K, or Alibi models like MPT. In contrast, Ms-PoE causes instability of the output of Qwen and Gemma in KV retrieval, leading to very low accuracy.
From the evaluation results of Longbench in Table 2, our method demonstrates varying degrees of improvement across different tasks, with the most significant increases being 1.5% in few-shot learning tasks, 3.4% in code tasks, 4% in synthetic tasks, 9.2% in single document QA tasks, and 1.9% in multi-document QA tasks. However, maybe because Longbench mainly focuses on comprehensiveness and reality, the influence of position bias on these tasks is relatively minimal, our method does not significantly improve the average scores (only a little). But it at least demonstrates that it will not impair the model’s original capability to handle various long context tasks.
@@ 4.3 Analysis
As shown in Table 1, there is an trend that our method generally increases performance when the key information is at the middle or end, but decreases when at the beginning. It reveals a possible fact that the positional hidden states may be an important factor causing the model to overly focus on the beginning parts of the context while miss the rear or middle parts. Therefore, scaling positional hidden states by a scaling factor less than 1 can reduce its impact, thus shift the model’s attention distribution from the bias to the beginning to a more balanced distribution.
In theory, a scaling factor over 1 amplifies the effect of positional hidden states, and one between 0 and 1 shrinks it. A negative factor can even reverse it. Correspondingly, as shown in Figure 6, a positive scaling factor over 1 causes the model to focus more on the beginning, while a negative factor shifts the focus towards the end. A factor between 0.5 and -1 leads to a relatively balanced attention distribution, where the average accuracy also peaks. These results demonstrate that scaling positional hidden states are essentially steering the position information in the model, to shift the position bias towards our desired direction.
@@
Table 3: Performance of difference models on MMLU and the timeline reorder task, before and after applying our method.
(a) w/ Attention
(b) w/ Performance
Figure 6: Attention distribution on KVs at different positions (depths) and performance when scaling the 2,393rd channel of Vicuna-7b with different scale factors on KV retrieval (Liu et al., 2024) of 100 KV pairs.
Side Effects
It is natural to concern scaling hidden states may impair the model’s normal workflow. So we utilized MMLU benchmark (Hendrycks et al., 2021), which assesses LLMs’ general capabilities, and the timeline-reorder task from LooGLE (Li et al., 2024), which requires arranging events chronologically across extensive text and is sensitive to positional information, to evaluate whether our method will impair the original abilities of LLMs. As shown in Table 3, there is no significant detriment to the models’ performance with our methods in default settings.
5 Conclusion
We propose a method which scales positional hidden states to mitigate the position bias issue of LLMs. Specifically, our study first explore the relationships between position bias (in the model’s performance), attention weights, causal mask, positional information in hidden states and positional channels, to confirm the positional channels can affect position bias. Based on these findings, we design a positional channel search algorithm to identify positional hidden states, and mitigate the model’s position bias by scaling the positional hidden states. These findings provide a new perspective for research related to position information and position bias of LLMs.
Limitations
The selection of layers to apply is mainly based on engineering experience. And although our method has tried to minimize the side effect of scaling hidden states, if choosing a too large scale factor or an inappropriate channel to scale, we still observe significant performance degradation.
Some newly released models like Qwen2.5 Qwen et al. (2024) have achieved nearly perfect performance on retrieval-based long-context tasks. Thus, on traditional benchmarks, our method may not bring more improvement. Experiment on more newly released models and benchmarks may still be needed in the future.
Appendix A Related Works
@@
Recent works reveal that LLMs often exhibit position bias, also known as the "lost in the middle" phenomenon (Liu et al., 2024; Kamradt, 2023). Previous efforts to mitigate this bias fall into several categories: 1) RoPE-based methods: These approaches modify the RoPE computation process to alleviate long-distance attention decay (less attention means less information retrieved), including Attention Bucket (Chen et al., 2024), which uses an ensemble of multiple RoPE bases to mitigate position bias, and Ms-PoE (Zhang et al., 2024), which dynamically interpolates with a small coefficient for different heads. 2) SFT-based methods (He et al., 2024a; Yu, 2023; An et al., 2024): These methods construct data with more diverse key information distributions or employ system2think SFT tasks to mitigate position bias. They require further training of the model. 3) Attention mask-based methods (He et al., 2024b): These methods modify attention mechanisms, including Attention Transition (Gao et al., 2023), which redirects attention to significant parts of the context and Stable Mask (Yin et al., 2024), which introduces pseudo attention into the causal mask, ensuring stable attention distribution when facing lengthy texts. 4) Prompt-based methods (Jiang et al., 2024; Peysakhovich and Lerer, 2023): These methods introduce an external module to reorder or compress information in the prompt, thereby mitigating position bias.
Appendix B Datasets Details